# Prompt Mixing in Diffusion Models using the Black Scholes Algorithm

### v1

@@ Abstract
We introduce a novel approach for prompt mixing, aiming to generate images at the intersection of multiple text prompts using pre-trained text-to-image diffusion models. At each time step during diffusion denoising, our algorithm forecasts predictions w.r.t. the generated image and makes informed text conditioning decisions. To do so, we leverage the connection between diffusion models (rooted in non-equilibrium thermodynamics) and the Black-Scholes model for pricing options in Finance, and draw analogies between the variables in both contexts to derive an appropriate algorithm for prompt mixing using the Black Scholes model. Specifically, the parallels between diffusion models and the Black-Scholes model enable us to leverage properties related to the dynamics of the Markovian model derived in the Black-Scholes algorithm. Our prompt-mixing algorithm is data-efficient, meaning it does not need additional training. Furthermore, it operates without human intervention or hyperparameter tuning. We highlight the benefits of our approach by comparing it qualitatively and quantitatively to other prompt mixing techniques, including linear interpolation, alternating prompts, step-wise prompt switching, and CLIP-guided prompt selection across various scenarios such as single object per text prompt, multiple objects per text prompt and objects against backgrounds. Code is available at https://github.com/divyakraman/BlackScholesDiffusion2024.
@@ 1 Introduction
Text-to-image diffusion models [13, 14] such as Imagen [33] by Google, Firefly [2] by Adobe, DALLE [7] by Microsoft, and Stable Diffusion [31, 11] by Stability AI can produce photorealistic images from textual descriptions and are typically trained on vast amounts of data using substantial computational resources. The latest versions of these models (as of May 2024) have seen significant improvements in comparison to their initial versions, attributed to various factors, such as increased training data, enhanced text understanding networks, improved image generation networks, and optimized strategies. Data efficient solutions [36], leveraging pre-trained diffusion models, can enhance results without requiring additional data collection. Prompt engineering [37] is one effective approach, involving techniques such as rephrasing prompts to improve model generalization. Another method is using large language models (LLMs) [23, 38] to parse complex prompts and determine useful priors for image generation. Attention maps [9] and prompt mixing techniques [27, 5] also contribute to better results.
Prompt mixing becomes particularly valuable when aiming to generate images that blend different text-image concepts i.e. when aiming to generate images at the intersection [27, 30, 28] of different text-image concepts. For instance, consider generating an image resembling a hybrid of a pink cat and a dog. Current diffusion models [2, 33, 7, 11] excel at straightforwardly combining individual text prompts to create visually appealing combinations. However, these capabilities are often restricted to advanced model versions (version 3) and rely on large-data techniques.
Data efficient prompt mixing techniques can enhance image generation without additional data. While linear interpolation [17] of text embeddings from individual prompts is a simple approach, it may not be optimal due to the highly non-linear nature of the text-image manifold and potential bias issues. Alternatives include switching between prompts during diffusion denoising, either by alternating [18] or using a step-wise technique [27]. However, these methods often require human involvement and careful prompt engineering for optimal results.
The question arises: what’s the best way to switch between text prompts for prompt-mixing? An automated approach considers the model’s varying capabilities with different prompts. While attention maps [15, 9] and layout guidance![42, 10] methods excel at guiding the model toward distinct scene entities, they may not be as effective for blending concepts within the same entity. To address this, during each step of diffusion denoising, the network should automatically focus on the text prompt that corresponds to the image’s deficiencies.
@@
In this paper, we introduce a novel approach to prompt mixing by integrating concepts from economics and finance. Our objective is to leverage pretrained diffusion models for prompt-mixing to generate images at the intersection of various text-image manifolds. During each step of the diffusion denoising process, our method dynamically conditions on the text prompt that requires the highest level of attention. To achieve this, our algorithm assesses the ‘cost’ associated with each relevant text prompt and selects the optimal conditions for the subsequent step of image generation based on the ‘cost’ that requires maximum optimization.
Key Insight: The inspiration for our approach lies in non-equilibrium thermodynamics [35], which is at the core of the development of diffusion models. These models share a conceptual foundation with the Black Scholes model[26], a Nobel Prize-winning mathematical framework extensively employed in financial markets for pricing European call options[8]. The denoising steps within a diffusion model, aimed at image generation, constitute a Markovian time-series. In our analogy, the image itself represents a valuable ‘stock’ or ‘asset’ that we aim to ‘purchase’ (or generate) at the most favorable ‘cost’ (or alignment with the text prompts). During each step of the diffusion denoising process, we extrapolate the generate latents to compute the ‘stock’ prices at the corresponding time-step, and use the properties of diffusion models to compute the various variables involved in the Black Scholes algorithm. Consequently, we leverage the Black Scholes model to predict a score for each text prompt. This score serves as an indicator of how the image should be conditioned in the subsequent timestep. By adopting this approach, our model dynamically focuses on aspects that require attention, ultimately generating an image that optimally satisfies all relevant text prompts.
We perform experiments on prompts with varying complexities to assess our method’s ability to seamlessly blend different objects and backgrounds across scenarios. Through qualitative and quantitative comparisons against several baselines-such as linear interpolation, prompt switching, and CLIP score-we highlight the superiority of our approach.
@@ 2 Related work
Prompt mixing.
Prompt mixing is a technique where different text prompts are used at various steps during the diffusion denoising process. Patashnik et al. [27] follow a step-wise approach, while Aerial Diffusion [18] alternates between prompts to create semantically consistent aerial-view images. Tools such as Image Mixer Diffusion [28] and CLIP Guided Image Mixing [30] follow similar techniques of prompt switching to blend multiple text prompts. However, these methods often require complex hyperparameter tuning.
Mixing step.
The mixing time or mixing step [22] of a Markov chain refers to the duration it takes for the chain to reach its steady-state distribution. Mixing time has found applications in image editing [46] and synthesizing out-of-distribution (OOD) images [45]. When switching between prompts using mixing time, a step-wise approach is followed, akin to the work by Patashnik et al. [27]. The advantage of mixing-time approaches over the method proposed by Patashnik et al. [27] is that it mathematically determines the optimal switching time without requiring complex hyperparameter tuning. To approximate the mixing time, it estimates the radius of the latent space. However, there are a few limitations to these methods. Firstly, most of these approaches are applied to diffusion models trained on relatively small, problem-specific datasets. Secondly, identifying clear boundaries on large-scale foundation models remains an unsolved challenge. Additionally, these methods lack the flexibility to choose the most optimal prompt at each timestep.
Understanding the latent space.
Tangential to our work, there has been substantial work [16, 1, 12] on exploring the latent space. These investigations have yielded valuable insights and practical solutions for downstream tasks such as image editing and manipulation applications [44, 34, 21, 43, 29]. Furthermore, a deeper understanding of the latent space within diffusion models has led to advancements in various methods [31, 21, 40].
3 Background
@@ 3.1 The Black Scholes Model
In this section, we provide a brief overview of the Black Scholes pricing model [26] used to determine the price of European call options of assets. which is used to determine the price of European call options on assets. In simple terms, a European call option allows an investor to lock in the price of an asset at any time but permits stock purchase (if desired) only upon expiration. Regardless of whether the stock price moves favorably or unfavorably over time, this option structure remains consistent. Investors rely on the Black-Scholes model to predict stock prices over time and make informed decisions about the optimal timing for stock purchases. The Black Scholes formula involves $5$ key variables:
@@
Assumptions made by the Black Scholes Model.
The Black Scholes model makes five assumptions: (i) No dividends are paid out during the life of an option, (ii) Market movements are somewhat random, (iii) There are no transaction costs in buying the asset, (iv) The volatility and risk free rate of the underlying asset are known and constant, (v) The returns of the underlying asset are normally distributed.
@@ 3.2 Relation to Diffusion Models
The main concept behind diffusion models [13] involves iteratively adding small amounts of random Gaussian noise to transform an initial photorealistic image $x_{0}$ to noise $x_{T}\sim\mathcal{N}(0,I)$ in $T$ steps. This process is known as the forward process. The converse of the theorem states that by starting from random noise $x_{T}\sim\mathcal{N}(0,I)$ and refining it iteratively for $T$ steps, we can generate a photorealistic image $x_{0}$. Since diffusion is a gradual process, the value of $T$ is typically large. At each intermediate timestep, $t\in\{0....T\}$, $x_{t}$ satisfies $x_{t}=\sqrt{\alpha_{t}}x_{0}+\sqrt{1-\alpha_{t}}\epsilon_{t}$. $0=\alpha_{T}<\alpha_{T-1}....<\alpha_{1}<\alpha_{0}=1$ are the hyperparameters of the diffusion schedule [13] and $\epsilon_{T}\sim\mathcal{N}(0,I)$. To obtain $x_{t-1}$ at each refinement step, the neural network $f_{\theta}(x_{t},t)$ is applied along with the corresponding random Gaussian noise perturbation. In other words, during each step of diffusion denoising, the known variance in added noise follows a Gaussian distribution [13]. In finance, the volatility refers to the standard deviation in the way stock prices change from one time instant to the next. The mathematical description of diffusion models, derived from non-equilibrium thermodynamics  [35], shares similarities with the derivation of the Black-Scholes model [26] used in pricing European call options within financial markets. Both models emerge from similar assumptions and conditions, underpinned by a shared mathematical structure. Assumptions (i) and (iii) made by the Black Scholes model, which talk about dividends and transaction costs, are irrelevant to diffusion models.
@@
4 Prompt Mixing using Black Scholes Model
@@ 4.1 Problem Formulation
Our objective is to generate an image that aligns with multiple text prompts simultaneously i.e. at the intersection of the various text-image manifolds of the diffusion models. We pose this task as a prompt mixing [27] problem. During each step of diffusion denoising, our aim is to select the most relevant text prompt (or the text prompt with respect to which the image requires further refining) for conditioning the model, ensuring the generation of an optimal image satisfying all individual text prompts.
@@ 4.2 Method
Selecting the most effective text prompt to condition a text-to-image diffusion model can be approached in various ways. Naively switching [18, 27] between prompts is sub-optimal due to the significant human effort involved. Moreover, it fails to address the critical issue of prioritizing the most relevant prompt in an automatic manner. Alternatively, one can compute the CLIP score at each step and move toward the text prompt with the lowest CLIP score. While this approach is reasonable, it overlooks essential factors related to the dynamics [13] of the diffusion denoising process, which significantly impact the final image generation. Leveraging the Black Scholes model allows us to model these dynamics effectively.
We utilize the CLIP score to quantify the “price” of the generated image, treating it as analogous to a stock. The CLIP score has been widely used to measure text-image alignment, making it a suitable metric for evaluating how well our generated image aligns with each text prompt. Consider diffusion denoising at step $i$ out of total steps $T$. We derive the relevant variables for Black Scholes formula as follows:

### v2

# Financial Models in Generative Art: Black-Scholes-Inspired Concept Blending in Text-to-Image Diffusion

@@ Abstract
We introduce a novel approach for concept blending in pretrained text-to-image diffusion models, aiming to generate images at the intersection of multiple text prompts. At each time step during diffusion denoising, our algorithm forecasts predictions w.r.t. the generated image and makes informed text conditioning decisions. Central to our method is the unique analogy between diffusion models, which are rooted in non-equilibrium thermodynamics, and the Black-Scholes model for financial option pricing. By drawing parallels between key variables in both domains, we derive a robust algorithm for concept blending that capitalizes on the Markovian dynamics of the Black-Scholes framework. Our text-based concept blending algorithm is data-efficient, meaning it does not need additional training. Furthermore, it operates without human intervention or hyperparameter tuning. We highlight the benefits of our approach by comparing it qualitatively and quantitatively to other text based concept blending techniques, including linear interpolation, alternating prompts, step-wise prompt switching, and CLIP-guided prompt selection across various scenarios such as single object per text prompt, multiple objects per text prompt and objects against backgrounds. Our work shows that financially inspired techniques can enhance text-to-image concept blending in generative AI, paving the way for broader innovation. Code is available at https://github.com/divyakraman/BlackScholesDiffusion2024.
@@ 1 Introduction
Text-to-image diffusion models [13, 14] have become essential tools for creative applications such as content creation, editing, personalization, and zero-shot generation. However, a core challenge remains: generating new and unseen concepts without retraining the model or collecting additional task-specific data. One promising avenue is to leverage the model’s intrinsic understanding of various conceptual manifolds to synthesize novel outputs by operating at their intersection.
This process, known as concept blending or prompt mixing [26, 5, 29, 27], used interchangeably in this paper, involves integrating two distinct concepts to produce a single, coherent image. For example, one could generate an image that merges the features of a pink cat with those of a dog. Although state-of-the-art diffusion models [2, 32, 7, 11] can blend text prompts into visually appealing results, these capabilities typically depend on advanced architectures and large-scale training data.
In contrast, data-efficient prompt mixing techniques hold the potential to enhance image generation while avoiding extra data collection costs. While simple strategies like linear interpolation of text embeddings [17] exist, they may not be optimal given the highly nonlinear nature of the text-image manifold and the risks of bias. Other methods, such as alternating [18] or step-wise prompt switching [26], encounter challenges stemming from human involvement and meticulous prompt engineering.
The central question we address is: How can we optimally switch between text prompts for effective concept blending? An ideal solution would automatically account for the varying capabilities of the model with respect to different prompts. While attention maps [15, 9] and layout guidance [40, 10] excel at directing the model toward distinct scene entities, they are less effective for blending concepts within the same entity. To overcome this, we propose an automated mechanism in which, at each step of the diffusion denoising process, the model focuses on the text prompt most aligned with the current image deficiencies.
@@
In this paper, we introduce a novel approach for concept blending that incorporates ideas from economics and finance. Our method leverages pre-trained diffusion models to generate images that reside at the intersection of various text-image manifolds, dynamically conditioning the generation process on the prompt demanding the greatest attention. Specifically, our algorithm evaluates the “cost” associated with each text prompt and selects the optimal conditioning for the subsequent generation step based on which prompt requires the most optimization effort.
Key Insight: Our approach is inspired by non-equilibrium thermodynamics [34]—the foundational principle behind diffusion models—and shares a conceptual basis with the Black–Scholes model [25, 8], a celebrated framework in financial markets for pricing European stock options. In our analogy, the image functions as a valuable asset whose generation should occur at the most favorable “cost,” determined by its alignment with the input text prompts. Concretely, at each denoising step we:
  1. 1.
    Extrapolate the latent representations to compute analogous “stock prices” at each timestep.
  2. 2.
    Leverage structural properties of diffusion models to derive the variables in the Black–Scholes algorithm.
  3. 3.
    Predict a score for each text prompt, indicating how the image should be conditioned in subsequent steps.
This mechanism enables our model to adaptively concentrate on the aspects most in need of refinement, ultimately generating images that optimally reflect all relevant text inputs.
We validate our approach through extensive experiments on prompts of varying complexity, demonstrating the model’s ability to seamlessly blend different objects and backgrounds. Quantitative and qualitative comparisons against baselines such as vanilla Stable Diffusion, linear interpolation, alternating sampling, prompt switching, and CLIP-guided prompt switching confirm the superiority of our method over current state-of-the-art techniques. Our work demonstrates that integrating financial methodologies into text-to-image synthesis can significantly improve concept blending, laying the groundwork for broader innovations in generative AI.
@@ 2 Related work
2.1 Concept Blending / Prompt Mixing
Prompt engineering [35] employs techniques such as rephrasing prompts to enhance model generalization. Additionally, large language models (LLMs) [23, 36] have been used to parse complex prompts and extract useful priors for image generation. Complementary to these methods are techniques involving attention maps [9] and prompt mixing, which applies different text prompts at various steps during the diffusion denoising process. For instance, Patashnik et al. [26] adopt a step-wise approach, whereas Aerial Diffusion [18] alternates prompts to produce semantically consistent aerial-view images. Similar strategies are employed by tools such as Image Mixer Diffusion [27] and CLIP Guided Image Mixing [29]. A common drawback is that these methods often necessitate complex hyperparameter tuning.
2.2 Mixing Step
The mixing time (or mixing step) of a Markov chain, as discussed in [22], represents the time required for the chain to reach its steady-state distribution. This concept has been exploited in areas such as image editing [44] and the synthesis of out-of-distribution (OOD) images [43]. When using mixing time for prompt switching—similar to the step-wise methods of Patashnik et al. [26]—one key advantage is that the optimal switching time is determined mathematically, thereby reducing the need for extensive hyperparameter tuning. Typically, the mixing time is approximated by estimating the radius of the latent space. However, these approaches often apply to diffusion models trained on relatively small, problem-specific datasets and may struggle with the challenges of larger-scale foundation models. Moreover, they generally lack the flexibility to select the most optimal prompt at every timestep.
2.3 Understanding the Latent Space
Exploration of the latent space has been a significant area of investigation [16, 1, 12]. Insights from these studies have facilitated advancements in downstream tasks such as image editing and manipulation [42, 33, 21, 41, 28]. Furthermore, a deeper understanding of the latent space in diffusion models has spurred improvements in various methods [30, 21, 38], providing a foundation for more effective image synthesis and manipulation techniques.
3 The Black Scholes Algorithm and Diffusion Models
@@ 3.1 The Black Scholes Model
In this section, we provide a brief overview of the Black Scholes pricing model [25] used to determine the price of European call options of assets. In simple terms, a European call option allows an investor to lock in the price of an asset at any time but permits stock purchase (if desired) only upon expiration. Regardless of whether the stock price moves favorably or unfavorably over time, this option structure remains consistent. Investors rely on the Black-Scholes model to predict stock prices over time and make informed decisions about the optimal timing for stock purchases. The Black Scholes formula involves $5$ key variables:
@@ 3.2 Relation to Diffusion Models
3.2.1 Diffusion Models
The main concept behind diffusion models involves iteratively adding small amounts of random Gaussian noise to transform an initial photorealistic image $x_{0}$ into noise $x_{T}\sim\mathcal{N}(0,I)$ over $T$ steps. This process is known as the forward process. Conversely, starting from random noise $x_{T}\sim\mathcal{N}(0,I)$ and refining it iteratively for $T$ steps can generate a photorealistic image $x_{0}$. Since diffusion is gradual, $T$ is typically large. At each intermediate timestep $t\in\{0,\ldots,T\}$, $x_{t}$ satisfies:
$x_{t}=\sqrt{\alpha_{t}}x_{0}+\sqrt{1-\alpha_{t}}\epsilon_{t}$
The hyperparameters of the diffusion schedule are $10=\alpha_{T}<\alpha_{T-1}<\ldots<\alpha_{1}<\alpha_{0}=$ 1, and $\epsilon_{T}\sim\mathcal{N}(0,I)$. To obtain $x_{t-1}$ at each refinement step, the neural network $f_{\theta}(x_{t},t)$ is applied along with the corresponding random Gaussian noise perturbation. Essentially, during each step of diffusion denoising, the known variance in added noise follows a Gaussian distribution.
3.2.2 Analysis based on Thermodynamics
The mathematical formulation of diffusion models, which is rooted in non-equilibrium thermodynamics, exhibits striking similarities to the derivation of the Black-Scholes model used for pricing European call options in financial markets. Both models arise from analogous assumptions and share a common mathematical structure.
Diffusion Models
In the context of diffusion, tools from statistical mechanics reveal that the generative dynamics undergo phase transitions and symmetry breaking. The dynamic equation of the generative process can be interpreted as a stochastic adiabatic transformation that minimizes free energy while maintaining thermal equilibrium. Moreover, diffusion models incorporate Gaussian noise in a controlled manner to learn the underlying data distribution: forward process, driving the system out of equilibrium by adding noise and reverse process, guiding the system back to a structured state.
The Black-Scholes Model
The Black-Scholes model assumes that the stock price $S_{t}$ follows a geometric Brownian motion, described by a stochastic differential equation (SDE). With the application of Itô’s Lemma, one derives a partial differential equation (PDE) that characterizes the evolution of option prices. Here, the pricing of derivatives can be viewed as a process of minimizing a form of “financial free energy” under specific constraints, analogous to thermodynamic systems seeking equilibrium.
A Unified Thermodynamic View
Both models can be interpreted through the lens of thermodynamics:
  • In diffusion models, the reverse process minimizes a free energy functional:
[…]
    where $q(\mathbf{x})$ represents the target distribution.
  • In the Black-Scholes framework, option pricing is analogous to minimizing a financial free energy, where the equilibrium state corresponds to a fair, risk-neutral price.
Thus, while diffusion models and the Black-Scholes equation address different applications—one in generative modeling and the other in finance—they both embody the principles of free energy minimization and equilibrium. The forward process in diffusion models perturbs the system away from equilibrium using controlled noise, whereas the Black-Scholes approach interprets market fluctuations through Brownian motion, seeking an equilibrium price despite the inherent randomness in financial markets.
3.2.3 Analysis based on SDEs
Both diffusion models and the Black-Scholes equation rely on stochastic differential equations (SDEs) to describe the evolution of systems over time. This common mathematical framework allows us to draw interesting parallels between seemingly disparate domains like image generation and financial option pricing.
Diffusion Models
In text-to-image diffusion models, the forward diffusion process is governed by the SDE:
$\mathbf{x}_{t}=-\frac{1}{2}\beta_{t}\mathbf{x}_{t}\,dt+\sqrt{\beta_{t}}\,d\mathbf{w}_{t},$
where $\mathbf{x}_{t}$ is the state (i.e., the image representation) at time $t$, $\beta_{t}$ is a time-dependent noise coefficient, and $d\mathbf{w}_{t}$ denotes the increment of a Wiener process (Brownian motion).
To reconstruct the data, the model uses the reverse diffusion process:
$\mathbf{x}_{t}=\left(\frac{1}{2}\beta_{t}\mathbf{x}_{t}+\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})\right)dt+\sqrt{\beta_{t}}\,d\mathbf{w}_{t},$
where $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})$ is the score function, guiding the reverse process.
Black-Scholes Equation
In finance, the Black-Scholes framework models the stock price $S_{t}$ as following a geometric Brownian motion:
$dS_{t}=\mu S_{t}\,dt+\sigma S_{t}\,dW_{t},$
with $\mu$ as the drift rate, $\sigma$ as the volatility, and $dW_{t}$ as the Wiener process.
Using Itô’s Lemma, the evolution of the option price $V(S,t)$ is given by:
$dV=\left(\frac{\partial V}{\partial t}+\mu S\frac{\partial V}{\partial S}+\frac{1}{2}\sigma^{2}S^{2}\frac{\partial^{2}V}{\partial S^{2}}\right)dt+\sigma S\frac{\partial V}{\partial S}\,dW_{t}.$
This derivation leads to the well-known Black-Scholes partial differential equation (PDE):
$\frac{\partial V}{\partial t}+rS\frac{\partial V}{\partial S}+\frac{1}{2}\sigma^{2}S^{2}\frac{\partial^{2}V}{\partial S^{2}}-rV=0,$
where $r$ is the risk-free interest rate.
Connecting the Two Domains
Both approaches yield PDEs describing the evolution of key quantities:
  1. 1.
    In diffusion models, the Fokker-Planck equation governs the probability density $p(\mathbf{x},t)$ of the image states:
    $\frac{\partial p(\mathbf{x},t)}{\partial t}=-\nabla\cdot\left(\mathbf{f}(\mathbf{x},t)p(\mathbf{x},t)\right)+\frac{1}{2}\nabla\cdot\left(D(\mathbf{x},t)\nabla p(\mathbf{x},t)\right),$
    where $\mathbf{f}(\mathbf{x},t)$ is the drift term and $D(\mathbf{x},t)$ the diffusion coefficient.
  2. 2.
    In the Black-Scholes model, the option price $V(S,t)$ evolves according to a PDE derived from the stock’s stochastic dynamics.
Moreover, both models use gradients to guide their respective processes. In diffusion models, the score function $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})$ directs the reverse diffusion process, while in the Black-Scholes framework, the gradient of $V(S,t)$ with respect to $S$ (often referred to as delta) is central to constructing a risk-neutral portfolio.
This similarity highlights a fascinating cross-domain synergy where concepts from financial mathematics can inspire novel techniques in generative modeling, and vice versa.
@@ 3.2.3 Analysis based on SDEs
4 Blending Concepts using the Black Scholes Algorithm
@@ 4.1 Problem Formulation
Our objective is to generate an image that aligns with multiple text prompts simultaneously i.e. at the intersection of the various text-image manifolds of the diffusion models. We pose this task as a prompt mixing or concept blending [26] problem. During each step of diffusion denoising, our aim is to select the most relevant text prompt (or the text prompt with respect to which the image requires further refining) for conditioning the model, ensuring the generation of an optimal image satisfying all individual text prompts.
@@ 4.2 Method
Selecting the most effective text prompt to condition a text-to-image diffusion model can be approached in various ways. Naively switching [18, 26] between prompts is sub-optimal—not only does it require significant human intervention, but it also fails to automatically prioritize the most relevant prompt. An alternative approach computes the CLIP score at each step and moves toward the prompt with the lowest CLIP score; however, this method overlooks the dynamic aspects inherent in the diffusion denoising process [13]. To address these issues, we leverage the Black-Scholes model to effectively capture these dynamics.
4.2.1 Black-Scholes Analogy in Diffusion Models
Our approach uses the CLIP score to quantify the “price” of the generated image, treating it as the underlying asset in a stock pricing analogy. The CLIP score—which is widely used to assess text-image alignment—provides a natural metric for how well the image aligns with each text prompt. For diffusion denoising at step $i$ out of a total of $T$ steps, we derive the Black-Scholes variables as follows: